Find the equation in the form of y=mx+b through (2,2), perpendicular to y=3x+4.
Answer by lwsshak3(11628)

Find the equation in the form of y=mx+b through (2,2), perpendicular to y=3x+4.
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
given equation:y=3x+4
slope,m=3
slope of line perpendicular to given line=-1/3(negative recprocal)
Equation:y=-x/3+b
solve for b using coordinates of given point(2,2)
2=-2/3+b
b=2+2/3=8/3
equation of perpendicular line:
y=-x/3+8/3